OpenStudy (anonymous):

Solve: 12.9 = x + 7.1

OpenStudy (anonymous):

Subtract 7.1 from both sides... x=5.8

OpenStudy (anonymous):

12.9= x + 7.1 12.9 - 7.1= x + 7.1 - 7.1 x=5.8

OpenStudy (anonymous):

txs